OpenStudy (jolo_yolo):

4. In a biological lab, the cell growth rate of two different organisms is tracked and recorded each week. Given the growth rate, the number of organisms can be determined using the following equations: s(x) = 100 + 23x m(x) = 90(1.2x)

OpenStudy (jolo_yolo):

3. Graph the system of equations, and show the point of intersection. | 4. Explain what the points graphed for each line represent. | 5. Explain how you can determine the solution to the equation 100 + 23x = 90(1.2x) using the graph from part c. | 6. Find the point(s) of intersection, and explain what the intersection represents in the context of the problem. |

OpenStudy (aripotta):

ok do they give you a coordinate plane to graph them on?

OpenStudy (jolo_yolo):

no im allowed to use desmos tho

OpenStudy (aripotta):

ok i've never used desmos :s

OpenStudy (jolo_yolo):

it's easy, you insert the equation and it automatically graphs it and shows you point of intersection.

OpenStudy (aripotta):

oh well there we go for number 3 lol

OpenStudy (jolo_yolo):

am i supposed to graph s(x) = 100 + 23x nd m(x) = 90(1.2x)?

OpenStudy (aripotta):

yea

OpenStudy (jolo_yolo):

oh okay i can do that. how bout part 4?

OpenStudy (aripotta):

so the x of a point means the time and the y means the number of organisms. so what can say for number 4?

OpenStudy (jolo_yolo):

um I could probably say the the x points represents time and the y points represent the amount or number of organisms.

OpenStudy (jolo_yolo):

is that good?

OpenStudy (aripotta):

lol i'd say "each point on the lines represent the number of organisms after x amount time has passed"

OpenStudy (aripotta):

amount of*

OpenStudy (aripotta):

i have to go real quick. i'll try to be back asap. maybe 5 - 10 min

OpenStudy (jolo_yolo):

Okay

OpenStudy (aripotta):

alright i'm back

OpenStudy (jolo_yolo):

wb

OpenStudy (aripotta):

so where the lines intersect would be the solution

OpenStudy (jolo_yolo):

oh I see so that's for 5 (ty btw for ur help) how bout 6

OpenStudy (aripotta):

so remember in part 2, we found out that x = 1.8? substitute 1.8 for x in either equation and solve for y to find the point of intersection

OpenStudy (aripotta):

y = 100 + 23(1.8)

OpenStudy (jolo_yolo):

y = 141

OpenStudy (aripotta):

if we didn't round to 1.8 and kept the long decimal, we'd get 127

OpenStudy (aripotta):

so idk if they want 141 or 127

OpenStudy (jolo_yolo):

well im not sure bc the question actually was asking for an explanation - 6. Find the point(s) of intersection, and explain what the intersection represents in the context of the problem.

OpenStudy (aripotta):

well we first had to find the point of intersection which was what we were trying to do here :/ it's either (1.8, 127) or (1.8, 141)

OpenStudy (aripotta):

in context of the problem, the point of intersection means that after 1.8 time has passed, the two organisms have the same quantity, 127 (or 141)

OpenStudy (jolo_yolo):

i'll try to graph it on desmos and see if i get 127 or 141. what were the equations again

OpenStudy (jolo_yolo):

I think it'll be 127 tho

OpenStudy (aripotta):

i graphed it using geogebra and i get (1.8, 127). but if they were expecting you to do it algebraically, then it would be (1.8, 141)

OpenStudy (jolo_yolo):

Oh okay so for the explaining part, would I say that the intersection is the solution to the context problem? i think im lost now

OpenStudy (aripotta):

write this: "the point of intersection is (1.8, 127). in context of the problem, that means that after 1.8 time has passed, the two organisms have the same quantity, 127."

OpenStudy (jolo_yolo):

thank you so much for your help!

OpenStudy (aripotta):

\(\Large \boldsymbol{ \rlap{\color{#008AFF}{\hspace{2.5pt}n}\color{#0073FF}{\hspace{0.34pt}o} \hspace{3pt}\color{#005CFF}{p} \color{#0045FF}{\hspace{0.9pt}r} \color{#0017FF}{o} \color{#0200FF}{\hspace{1.7pt}b} \color{#3000FF}{\hspace{1.5pt}l} \color{#4700FF}{\hspace{0.4pt}e} \color{#7500FF}{\hspace{0.9pt}m} \color{#8C00FF}{\hspace{0.5pt}o}~\LARGE\color{#D100FF}{\ddot\smile}}{\hspace{1.5pt} no~p\hspace{0.3pt}r\hspace{0.4pt}o\hspace{1pt}b\hspace{1.5pt}l\hspace{0.4pt}e\hspace{0.4pt}m\hspace{0.4pt}o}}\)

OpenStudy (jolo_yolo):

nicee haha

OpenStudy (aripotta):

:3